OpenStudy (anonymous):

For what x-value(s) does sin(x) = -1?

OpenStudy (ikram002p):

well when \[x =\frac{ \pi }{ 2 } ,\sin\frac{ \pi }{ 2 } = 1 ,\] you need to know in which quarter the sins is negative that would be in 3 & 4 quarter , so \[x=\frac{ \pi }{ 2 }+n \pi \]

OpenStudy (austinl):

This may help!

OpenStudy (anonymous):

@ikram002p n = ?

OpenStudy (ikram002p):

n = any positive number you can use n= 1 or 2 or 3 or ....

OpenStudy (anonymous):

@ikram002p it won't let me see your answer anymore...

OpenStudy (anonymous):

pi/2 + n/pi ?

OpenStudy (ikram002p):

supose n = 1 \[x=\frac{ \pi }{ 2}+\pi =\frac{ 3\pi }{ 2 }\] now its better is that good for you ??